OpenStudy (anonymous):

Calculus: Find two numbers where the sum of the first number and four times the second number is 1000, and the product of both is as large as possible.

OpenStudy (anonymous):

I have noo idea how to even start that.

OpenStudy (amistre64):

let the numbers be a and b a + 4b = 1000 ab is a max

OpenStudy (amistre64):

define a or b in terms of the other and work out a quadratic to play with

OpenStudy (amistre64):

or you could play with langrange multipliers ...

OpenStudy (anonymous):

@amistre64 I got A as 500 and B as 125, does that sound right to you?

OpenStudy (amistre64):

maybe ... a + 4b = 1000; a = 1000-4b ab = (1000-4b)b = 1000b - 4b^2 the vertex is at: b = -1000/-2(4) = 125 so yes

OpenStudy (anonymous):

Thank you!